OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

hartnn (hartnn):

hint : 3x =2x+x , then 2x = x+x and use the formula \(\cos(A+B)=\cos A \cos B - \sin A\sin B\) twice, once when A=2x, B=x....then when A=x, B=x

OpenStudy (anonymous):

. . . =cosx( 2(cosx)^2 -1 -2(sinx)^2 ) ??

OpenStudy (anonymous):

then what would i get?

OpenStudy (anonymous):

cos(2x+x)=cos2xcosx-sin2xsinx; cos(x+x)=cosxcosx-sinxsinx=(cosx)^2 - (sinx)^2; sin(2x+x)=sin2xcosx+sinxcos2x; sin(x+x)=sinxcox+sinxcosx=2sinxcosx; cos3x=cos2xcosx-sin2xsinx; =((cosx)^2 - (sinx)^2)cosx - (2sinxcosx)sinx =( 2(cosx)^2 -1 )cosx -2cosx(sinx)^2 =2(cosx)^3-cosx-2cosx(sinx)^2 =(cosx) [ 2(cosx)^2 - 1 -2(sinx)^2) ] =(cosx) [ 4(cosx)^2 -3 ]

OpenStudy (anonymous):

It's either A. cosx -4 cosx sin^2x B. -sin^3x+2 sinx cosx C. -sin^2x+ 2sinx cosx D. 2sin^2xcosx-2sinxcosx

OpenStudy (anonymous):

A

OpenStudy (anonymous):

from my answer =(cosx) [ 4(cosx)^2 -3 ] =4(cosx)^3-3cosx substitute (cosx)^2 = 1 - (sinx)^2 into the question =4cosx(cosx)^2-3cosx =4cosx [ 1 - (sinx)^2 ]-3cosx =4cosx - 4cosx(sinx)^2-3cosx =cosx- 4cosx(sinx)^2.. A

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

your wc :D